1. Excelor stock is expected to pay $2.80 per share as its next annual dividend. The...

1.

Excelor stock is expected to pay $2.80 per share as its next annual dividend. The firm has a policy of increasing the dividend by 10.0 percent annually. The stock has a market price of $13.61 and a beta of 2.6. The market risk premium is 8.54 percent and the risk-free rate is 4.80 percent. What is the cost of equity?

27.71 percent

28.14 percent

29.18 percent

28.79 percent

2.

The 13.50 percent preferred stock of Ajax Unlimited is selling for $93.30 a share. What is the cost of preferred stock if the risk-free rate is 4.45 percent?

17.66 percent

14.47 percent

18.27 percent

13.53 percent
Answer #1

1. D. 28.79 percent

Dividend growth: ($2.80 / $13.61) + 0.10 = 0.3057

CAPM: 0.0480 + 2.6(0.0854) = 0.2700

Cost of equity = (0.3057 + 0.2700) / 2 = 28.79 percent

2. B. 14.47 percent

Cost of preferred = (0.1350 × $100) / $93.30 = 14.47 percent